OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. G(4,6) H(1,5)

OpenStudy (lgbasallote):

use \[y - y_1 - \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\]

OpenStudy (anonymous):

but can you please do it for me i dont understand it

OpenStudy (apoorvk):

Okay dokey!!! \[\text{For a line passing through two points (}x_1 y_1/text{) and (}x_2, y_2\text{), its equation is:} \]\[\frac{y_2 - y_1}{x_2 - x_1} = \frac{y_2 - y}{x_2 - x}\] Put in the values from the coordinates of the points given to you, and then rearrange in standard form.

OpenStudy (anonymous):

can you please do it

OpenStudy (apoorvk):

nada. i can help though. (x1, y1) = (4,6) (x2,y2) = (1,5) Just plug in the values!

OpenStudy (anonymous):

ok did it whats next

OpenStudy (apoorvk):

what equation did you get by doing that? Use the 'equation' editor, or the use the 'draw' option below.

OpenStudy (anonymous):

i got 24x divided by 5x

OpenStudy (apoorvk):

You aren't really trying are you? Just typing some random stuff to make us think you are, so that we may give you the answer. Am sorry, we are ready help people, but not to serve answers on a plate. You just need to stay calm, and read what I say above. This is as easy as spelling C-A-T cat, you juts have to follow what we are sayin'! :)